# Neural Guided Diffusion Bridges

### v2

@@ 5.1 Linear processes
We examine one-dimensional linear processes with tractable conditional drifts, such as a Brownian bridge with constant drift and the Ornstein-Uhlenbeck process. For these simple models, the lower bound of $L(\theta)$ can be analytically computed, providing a benchmark to verify whether the neural network achieves this bound. More details can be found in Appendix B.2.
Brownian bridge: Consider a one-dimensional diffusion with constant drift $\gamma$ and constant diffusion coefficient $\sigma$. As the transition density $p(t,x_{t}\mid s,x_{s})$ is Gaussian, the SDE for the process that is conditioned to hit $v$ at time $T$ is given by
@@
Suppose we construct the guided proposal using an auxiliary process $\tilde{X}$ with zero drift (i.e.  $\beta(t)=B(t)=0$) and diffusion coefficient $\tilde{\sigma}=\sigma$. The corresponding neural guided proposal $X^{\bullet}$ solves the SDE
@@
By comparing $X^{\bullet}$ with $X^{\star}$, it is clear that the optimal map $\vartheta_{\theta}$ is given by the map $\vartheta_{\theta_{\text{opt}}}$ defined by $\vartheta_{\theta_{\text{opt}}}(t,x)=-\gamma/\sigma$. Additionally, the lower bound of $L(\theta)$, $\log\frac{h(0,x_{0})}{\tilde{h}(0,x_{0})}$, is analytically tractable since the transition densities $\tilde{p}$ of $\tilde{X}$ are Gaussian.
Suppose $\gamma=\sigma=1.0$ and $x_{0}=v=0$. Figure 1 shows the empirical marginal distributions of $X^{\star}$ and $X^{\bullet}$ derived from independent samples from approximations of Equation (22) and Equation (23). In the topleft plot of Figure 9 we show how the training varies over iterations. The horizontal red line is the known lower bound on the training loss. In the bottomleft plot, we show a heatmap of the difference between the learned neural network and $\vartheta_{\theta_{\text{opt}}}$. The experiment has been repeated for other values of $\gamma$ ($2$, $5$ and $10$).
Figure 1: Comparison of marginal distributions at different time slices of the learned neural bridge and analytical true bridge. Top: Brownian bridges; Bottom: Ornstein-Uhlenbeck bridge, the pink bars show the error of each bin. The histograms are made from 1,000 independent trajectory samples individually.
We conclude that the neural net is able to learn the optimal drift and that the neural guided bridge is very close to the true bridge in terms of Kullback-Leibler divergence.
Ornstein-Uhlenbeck bridge: Consider the Ornstein-Uhlenbeck (OU) process
$\mathrm{d}X_{t}=-\gamma X_{t}\mathrm{d}t+\sigma\mathrm{d}W_{t}.$ (24)
@@
Upon choosing $\tilde{X}_{t}=\sigma W_{t}$, the neural guided bridge satisfies the SDE
@@
$\mathrm{d}X^{\bullet}_{t}=\left\{-\gamma X^{\bullet}_{t}+\frac{v-X^{\bullet}_{t}}{T-t}+\sigma\vartheta_{\theta}(t,X^{\bullet}_{t})\right\}\mathrm{d}t+\sigma\mathrm{d}W_{t}.$ (25)
In Appendix B.2 we derive the optimal map $\vartheta_{\theta_{\text{opt}}}$ and lower bound on $L$. With SDE parameters set as $x_{0}=v=0$, $\sigma=1.0$, we repeat the experiments of evaluating training loss curves and outputs of the learned neural network, as were done for the Brownian bridge and show the results in Figure 10.
A comparison of the results from Brownian and OU bridges reveals significant fluctuations in the training loss curves, particularly pronounced in the OU bridges. These fluctuations arise from the Monte Carlo approximation of the expectation and the numerical discretization of the integral in the training loss. The Monte Carlo approximation is implemented by computing the empirical mean over $N$ Wiener sample paths, while the integral is discretized on a finite grid $\mathcal{T}$ using left-point Riemann sums. The sampling introduces noise into the optimization, resulting in strong fluctuations. The discretization may cause bias in attaining the lower bound. In the OU case, $\vartheta_{\theta}$ exhibits a more complex form compared to the Brownian case, necessitating a finer grid and a larger number of Monte Carlo samples.
Despite the fluctuations observed during the training of the neural bridge, its validity under appropriate conditions is confirmed by quantitative evaluations in linear process experiments. For experiments lacking analytical references, performance is assessed qualitatively
@@ 5.2 Cell diffusion model
Wang et al. [2011] introduced a model for cell differentiation which serves as a test case for diffusion bridge simulation in Heng et al. [2022], Baker et al. [2024b]. Cellular expression levels $X_{t}=(X_{t,1},X_{t,2})$ are governed by the 2-dimensional SDE
@@
driven by a 2-dimensional Wiener process $W$. The highly nonlinear drift makes this a challenging case. The guided neural bridge is constructed by taking the auxiliary process that drops the nonlinearities in the drift. More precisely,
$\mathrm{d}\tilde{X}_{t}=-\tilde{X}_{t}\mathrm{d}t+\sigma\mathrm{d}W_{t}.$ (27)
We study three representative cases that result in distinct dynamics for the conditional processes: (1) events that are likely under the forward process, which we refer to as “normal” events; (2) rare events; and (3) events that cause trajectories to exhibit multiple modes, where the marginal probability at certain times is multimodal. We compare our method to (a) the guided proposal [Schauer et al., 2017]; (b) bridge simulation via score matching [Heng et al., 2022]; and (c) bridge simulation using adjoint processes [Baker et al., 2024b]. Details of the experiments are presented in Appendix B.3.
@@
Normal event: We set $x_{0}=[0.1,-0.1]^{T}$ and $v=[2.0,0.1]^{T}$. From multiple forward simulations, it can be assessed that balls around $v$ get non-negligible mass. We take $T=4.0$ and $\sigma=0.1$. In Figure 2, we show 30 conditional sample paths obtained by the three baseline methods mentioned above and the trained neural bridge. Since the true conditional process is analytically intractable, we sample 100,000 times from the forward (unconditional) process Equation (26) and retain samples that satisfy $\|LX_{T}-v\|\leq 0.01$. The obtained samples can be treated as a reference template for the true conditioned dynamics. Overall, all four methods successfully recover the truth dynamics. The performance of all four methods considered is comparable. We only note that the adjoint bridge sample paths appear slightly more dispersed.
Figure 2: Visualization of 30 simulated bridges when conditioning on normal events using different methods. Note that except for the guided proposal, all the samples are independently drawn. The underlying grey trajectories presented in the figure are the unconditional samples by forward solving $X$ without conditioning and filtered by $\|LX_{T}-v\|<0.01$ as the reference of the true dynamics.
Rare event: We set $x_{0}=[0.1,-0.1]^{T}$, $v=[1.0,-0.1]^{T}$, which is a rare event for $X$ to hit. In this experiment, we set $\sigma=0.1$. Unlike the normal event case, the true dynamics cannot be recovered by brute-force sampling from the unconditioned process, as it is highly improbable that paths end up in a small ball around $v$. We display 30 conditional paths each generated by the four methods considered in Figure 3. Visual similarity can be observed between the neural guided bridge and the guided proposal, while the adjoint bridge cannot effectively model reasonable dynamics, as also reported in Baker et al. [2024b]. Score matching performs worst in this setting, which is expected because the score is learned from unconditional samples which as time $T$ tend to be far from $v$. This scarcity leads to poor estimation in those areas, resulting in significant deviations in the initial stages.
Figure 3: Visualization of 30 simulated bridges when conditioning on rare events using different methods. Note that except for the guided proposal, all the samples are independently drawn. No reference trajectory is available when conditioned on rare events.
Multi-modality: In both of the previous cases, each component’s marginal distribution at times $(0,T]$ is unimodal. However, with some special initial conditions, multimodality can arise, which poses a challenging task where one would like to recover all modes. Specifically, let $x_{0}=[-1.0,-1.0]^{T}$, $v=[-1.0,-1.0]^{T}$, $T=5.0$ and $\sigma=0.1$. Figure 4 compares the performance of the different methods. Both the adjoint bridge and score matching fail to model the dynamics accurately. The neural guided bridge and the guided proposal have comparable performance, as they both can only cover part of the modes. Moreover, when running pCN for a single chain, sometimes iterates of samples get stuck at a single mode. To address this, multiple chains must be run, which may become computationally expensive. In contrast, once the neural bridge is trained, new samples can be immediately generated by sampling paths from the learned SDE, at a cost comparable to unconditional forward simulation, while still covering the majority of modes.
Figure 4: Visualization of 30 simulated bridges when conditioning on events that will cause multi-modality using different methods. Note that except for the guided proposal, all the samples are independently drawn. No reference trajectory is available when conditioned on multi-modality events.
Among all three tasks, our method demonstrates strong flexibility and adaptability, while the alternative methods show limitations when applied to specific tasks.
@@ 5.3 FitzHugh-Nagumo model
We consider a FitzHugh-Nagumo model, which is a prototype of an excitable system, considered for example in Ditlevsen and Samson [2019], Bierkens et al. [2020]. It is described by the SDE
@@
We condition the process by the value of its first component by setting $L=\begin{bmatrix}1&0\end{bmatrix}$ and hence condition on the event $LX_{T}=v$. We construct the guided proposal just as proposed in Bierkens et al. [2020] using the Taylor expansion $-x^{3}\approx 2v^{3}-3v^{2}x$ at $x=v$. Accordingly, we choose $\tilde{X}$
@@
Suppose $[\chi,s,\gamma,\alpha,\sigma]=[0.1,0,1.5,0.8,0.3]$. We examine the conditional behaviour of $X$ within $t\in[0,2.0]$ under two scenarios: (1) conditioning on a normal event $v=[-1.0]$; and (2) conditioning on a rare event $v=[1.1]$. As the score-matching and adjoint-process methods have not been proposed in the setting of a partially observed state, we only compare our method to the guided proposal. The results are shown in Figure 5 and Figure 6 respectively.
Normal event: Both the neural guided bridge and the guided proposal recover most patterns of the true process as shown in Figure 5. However, the neural guided bridge performs slightly better in capturing a broader spanning region of $X_{t,2}$ for $t\in(0.5,1.5)$. Additionally, $X_{t,1}$ generated by the neural guided bridge exhibits better visual consistency with the reference compared to the guided proposal.
Figure 5: Visualization of FHN model conditioned on a normal event using 30 samples each either by drawing independently from the neural guided bridge or subsampling from the guided proposal updating history, the reference is obtained by directly sampling original Equation (28), and filtered by the condition $\|LX_{T}-v\|\leq 0.01$.
Rare event: The conditioning process introduces multimodality into the bridge, causing both methods to occasionally sample from only a single mode, as anticipated based on the previous example.
The left-hand panel of Figure 6 shows “true” conditioned paths obtained by forward sampling and keeping that paths that satisfy a relaxed version of the conditioning. Clearly, the bridge process is bimodal. The neural guided bridge samples paths from only one of the two modes, though the sampled paths appear very similar to the actual conditioned paths. The guided proposal samples paths inbetween the two modes as well, which are not seen in the reference panel. These observations suggest that both methods struggle to recover the 2 modes.
Figure 6: Visualization of FHN model conditioned on a rare event using 30 samples each either by drawing independently from the neural guided bridge or subsampling from the guided proposal updating history, the reference is obtained by directly sampling original Equation (28), and filtered by the condition $\|LX_{T}-v\|\leq 0.01$.
@@ 5.4 Stochastic landmark matching
At last, we consider a relatively high-dimensional stochastic nonlinear conditioning task: stochastic landmark matching, mainly specified in [Arnaudon et al., 2022], where a geometric shape is discretized and represented as $X_{t}=\{X^{(i)}_{t}\}_{i=1,\dots,n}\subset\mathbb{R}^{nd}$ a finite set of $n$ distinct landmarks $X^{(i)}_{t}\in D\subseteq\mathbb{R}^{d}$ and $d=\{2,3\}$ generally indicates a 2 or 3-dimensional shape. The stochastic landmark model defined in [Arnaudon et al., 2022] without momentum (i.e., only the coordinate variable dynamics) is defined by an SDE with a set of $n$ $d$-dimensional Wiener processes $W_{t}=\{W_{t}^{(j)}\}_{j=1,\dots,n}\in\mathbb{R}^{nd}$:
$\mathrm{d}X_{t}=Q(X_{t})\mathrm{d}W_{t},\quad Q(X_{t})_{ij}\coloneq k(X_{t}^{(i)},X_{t}^{(j)}).$ (30)
$k$ is a kernel function $\mathbb{R}^{d}\times\mathbb{R}^{d}\to\mathbb{R}$. We choose a Gaussian kernel $k(x,y)=\alpha\exp\left(-\frac{\|x-y\|^{2}}{2\kappa^{2}}\right)$. Note that the diffusion coefficient $Q$ is state-dependent, which is necessary for maintaining the structure of the shape during the evolution. Figure 13 shows an intuitive comparison between using a linear process and Equation (30) to model the stochastic evolutions of the same shape. In modelling real-world stochastic shapes like medical anatomical markers or biological species outlines, such a structure-preserving property of Equation 30 is preferred.
We choose
$\mathrm{d}\tilde{X}_{t}=Q(X_{T})\mathrm{d}W_{t}.$ (31)
That is, we only evaluate the diffusion coefficient at the endpoint and keep it constant throughout the process. We choose one ellipse as the starting point and another ellipse as the endpoint of the bridge. Each ellipse is discretized as 50 landmarks, leading to the dimension of $X_{t}$ being 100. The kernel parameters are chosen to be $\alpha=0.3,\kappa=0.5$ to ensure a strong correlation between a wide range of landmarks. Let $T=1.0$. Figure 7 and Figure 8 show 4 samples of the learned neural guided bridge and the guided proposal. Both methods capture the correct dynamics and approach the target shape. Despite the visual comparison, there is no quantitative way of comparing the performance of the two methods as the true bridge is anyway intractable. However, we note that for such a high-dimensional problem, using the Metropolis-Hasting method to sample from the guided proposal is much slower than numerical solving the trained neural guided bridge, since the latter only requires a single forward solving, while the former requires computationally expensive MCMC update steps, especially when the dimension of the problem is high. In this case, training a neural bridge once and subsequently generating independent samples may be preferable.
Figure 7: Four independent samples of the stochastic bridges between two ellipses obtained by solving the learned neural diffusion bridge. The target ellipse is marked in yellow, and the colour of the trajectories indicates time.
Figure 8: Four samples of the stochastic bridges between two ellipses obtained by subsampling from the guided proposal updating history. The target ellipse is marked in yellow, and the colour of the trajectories indicates time.

### v3

@@ 5.1 Linear Processes
We consider one-dimensional linear processes with analytically tractable conditional drifts, including Brownian motion with constant drift and the Ornstein–Uhlenbeck process. For these models, the lower bound of $\theta\mapsto L(\theta)$ can be explicitly computed, serving as a benchmark to assess whether the neural network reaches this bound. Additional details are provided in Section B.2.
Brownian bridge: Consider a one-dimensional Brownian motion with constant drift: $\mathrm{d}X_{t}=\gamma\mathrm{d}t+\sigma\mathrm{d}W_{t}$. As its transition density $p(t,x_{t}\mid s,x_{s})$ is Gaussian, the fully-observed process conditioned on $\{X_{T}=v\}$, satisfies the SDE:
@@
We construct the guided proposal using the auxiliary process $\tilde{X}_{t}=\sigma W_{t}$. It is easy to see that $X^{\bullet}$ solves the SDE:
@@
By comparing $X^{\bullet}$ with $X^{\star}$, it is clear that the optimal map $\vartheta$ is given by $\vartheta_{\theta_{\text{opt}}}(t,x)=-\gamma/\sigma$. Additionally, the lower bound on $L$, $\log\frac{\tilde{h}(0,x_{0})}{h(0,x_{0})}$, is analytically tractable since the transition densities $\tilde{p}$ of $\tilde{X}$ are Gaussian. In Figure 6(a), we track how the training varies over iterations under different settings of $\gamma$ and $\sigma$, which leads to different lower bounds. It can be seen that all the trainings converge to corresponding theoretical lower bounds. Figure 2 compares the trained map $\vartheta_{\theta}$ with the optimal map $\vartheta_{\theta_{\text{opt}}}$. The neural network matches the optimal map in regions well supported by the training data, but the approximation error grows outside these regions—a common limitation of neural network training. Figure 3 shows the empirical marginal densities of the neural bridge alongside the analytical densities obtained from independent simulations of Equations 24 and 25. The learned and analytical distributions are in close agreement.
Ornstein-Uhlenbeck bridge: We now consider the Ornstein-Uhlenbeck (OU) process:
$\mathrm{d}X_{t}=\gamma(\mu-X_{t})\mathrm{d}t+\sigma\mathrm{d}W_{t},$ (26)
@@
which requires a $t,x$-dependent $\vartheta_{\theta_{\text{opt}}}$ to correct the guided proposal. Upon choosing $\tilde{X}_{t}=\sigma W_{t}$, the neural bridge satisfies the SDE:
@@
$\mathrm{d}X^{\bullet}_{t}=\left\{\gamma(\mu-X^{\bullet}_{t})+\frac{v-X^{\bullet}_{t}}{T-t}+\sigma\vartheta_{\theta}(t,X^{\bullet}_{t})\right\}\mathrm{d}t+\sigma\mathrm{d}W_{t}.$ (27)
As with Brownian motion, the OU process has Gaussian transition densities and is therefore analytically tractable. Using this property, we derive the optimal map $\vartheta_{\theta_{\text{opt}}}$ and the corresponding lower bound on $L$, given in Equations 47 and 46. We vary the parameters $\gamma$, $\mu$, and $\sigma$ and plot the resulting training-loss curves alongside the analytical lower bound in Figure 6(b); in every case the loss converges to its respective bound. Figure 4 compares the outputs of the trained network with the optimal map, and Figure 5 shows the empirical marginal densities of the learned neural bridge with those of the analytical bridge obtained from independent simulations of Equations 27 and 44.
From the above two experiments, we conclude that the neural network is able to learn the optimal drift with the proposed loss function and that the neural guided bridge is very close to the true bridge in terms of KL divergence (reflected by the very small differences between training losses and analytical lower bounds). In the remaining numerical examples no closed form expression is available for $h(0,x_{0})$ and therefore performance can only be assessed qualitatively.
@@ 5.2 Cell Diffusion Model
[Wang et al., 2011] introduced a model for cell differentiation which serves as a test case for diffusion bridge simulation in [Heng et al., 2022, Baker et al., 2024b]. Cellular expression levels $X_{t}=\begin{bmatrix}X_{t,1}&X_{t,2}\end{bmatrix}^{\top}$ are governed by the 2-dimensional SDE:
@@
driven by a 2-dimensional Wiener process $W$. The highly nonlinear drift makes this a challenging case. The guided neural bridge is constructed as an OU process:
$\mathrm{d}\tilde{X}_{t}=\{\mathbf{1}-\tilde{X}_{t}\}\mathrm{d}t+\sigma\mathrm{d}W_{t},$ (29)
[…]
(a) Unconditioned
(b) Normal event
(c) Rare event
(d) Multi-modality
Figure 1: Top left (a): $30$ realisations from the unconditioned cell diffusion model (Cf. Equation 28). Top right (b): $30$ realisations from the learned conditional process on a “normal event” $v=[2.0\,-0.1]^{\top}$; Bottom left (c): $30$ realisations from the learned conditional process on a rare event $v=[1.0\,-0.1]^{\top}$; Bottom right (d): $30$ realisations from the learned conditional process on event $v=[-1.0\,-1.0]^{\top}$ that causes multi-modality. Except in panel (d), all processes start from $x_{0}=[0.1\,-0.1]^{\top}$ and run up to time $T=4.0$. In (d), the process starts from $x_{0}=[-1.0\,-1.0]^{\top}$ and runs up to time $T=5.0$.
@@
Normal event: We set $x_{0}=\begin{bmatrix}0.1&-0.1\end{bmatrix}^{\top}$, $T=4$ and $v=\begin{bmatrix}2.0&-0.1\end{bmatrix}^{\top}$. From the topleft panel in Figure 1 it is clear that this corresponds to a normal event: balls around $v$ at time $T$ get non-negligible mass under the unconditioned process. In Figure 1(b), we show 30 sample paths obtained from sampling the trained neural bridge and Figure 7 shows a comparison to the three baseline methods mentioned above. Since the true conditional process is analytically intractable, we generated $100,000$ samples from the forward (unconditional) process Equation 28, and obtained $172$ samples that satisfy $\|LX_{T}-v\|\leq 0.01$, and only show first $30$ samples in the figure. Those samples can be treated as samples close to true bridges. Overall, all four methods successfully recover the true dynamics. The performance of all four methods considered is comparable.Note that the adjoint bridge sample paths appear slightly more dispersed.
Rare event: We set $x_{0}=\begin{bmatrix}0.1&-0.1\end{bmatrix}^{\top}$ and $T=4$ as before but now we take $v=\begin{bmatrix}1.0&-0.1\end{bmatrix}^{\top}$, which is a rare event. Unlike the “normal event” case, the true dynamics cannot be recovered by forward sampling from the unconditioned process, as it is highly improbable that paths end up in a small ball around $v$. In Figure 1(c), we show 30 sample paths obtained from sampling the trained neural bridge and Figure 7 shows a comparison to the three baseline methods mentioned above.
All methods except the adjoint forward approach capture the correct trajectory dynamics, consistent with findings in [Baker et al., 2024b]. Among the remaining three, trajectories from the neural bridge align more closely with those from the guided proposal than those from score matching. In the score matching method, samples of $X_{t,1}$ show higher variance prior to $t=3.5$, suggesting less accurate score estimates. This is expected, as score matching learns its drift from samples of the unconditioned process which rarely reaches small balls around $v$, leading to poor approximation quality.
Multi-modality: In both of the previous cases, each component’s marginal distribution at times $(0,T]$ is unimodal. However, with some special initial conditions, multimodality can arise, which poses a challenging task where one would like to recover all modes. Specifically, let $x_{0}=\begin{bmatrix}-1.0&-1.0\end{bmatrix}^{\top}$, $T=5.0$ and $v=\begin{bmatrix}-1.0&-1.0\end{bmatrix}^{\top}$. In Figure 1(d), we show 30 sample paths obtained from sampling the trained neural bridge and Figure 7 shows a comparison to the three baseline methods mentioned above.
Both the adjoint bridge and score matching fail to model the dynamics accurately. In contrast, the neural bridge and guided proposal yield similar marginal distributions. However, good performance of the guided proposal may take many MCMC iterations, or possibly the use of multiple (interacting) chains. Once trained, the neural bridge can generate independent samples, at a cost comparable to unconditioned forward simulations, while maintaining sampling quality close to that of the guided proposal.
Across all three tasks, the neural bridge shows strong adaptability and achieves performance comparable to the guided proposal, with the added benefit of faster independent sampling. In contrast, the other two methods exhibit limitations under specific settings.
@@ 5.3 FitzHugh-Nagumo Model
We consider the FitzHugh-Nagumo model, which is a prototype of an excitable system, considered for example in Ditlevsen and Samson [2019], Bierkens et al. [2020]. It is described by the SDE:
@@
We condition the process by the value of its first component by setting $L=\begin{bmatrix}1&0\end{bmatrix}$ and hence condition on the event $LX_{T}=v\in\mathbb{R}$. We construct the guided proposal as proposed in Bierkens et al. [2020] using the Taylor expansion $-x^{3}\approx 2v^{3}-3v^{2}x$ at $x=v$. Thus, $\tilde{X}$ satisfies the SDE
@@
Suppose $\{\chi,s,\gamma,\alpha,\sigma\}=\{0.1,0,1.5,0.8,0.3\}$. We examine the conditional behaviour of $X$ within $t\in[0,2.0]$ under two scenarios: (1) conditioning on a normal event $v=-1.0$; and (2) conditioning on a rare event $v=1.1$. As the score-matching and adjoint-process methods have not been proposed in the setting of a partially observed state, we only compare our method to the guided proposal.
Normal event: Figure 11 compares sample trajectories generated by the neural bridge and guided proposal. As a reference, paths obtained by unconditional sampling are added, where any path not ending close to the endpoint has been rejected. Both methods capture the key features of the conditioned dynamics, closely matching the reference trajectories. However, the neural bridge achieves this using independently drawn samples after training, whereas the guided proposal requires a long MCMC chain. The similarity between the neural bridge and the guided proposal confirms that the neural bridge effectively approximates the conditioned distribution, indicating accurate modeling of the dynamics under normal event conditioning.
Rare event: InFigure 12 we redo the experiment under conditioning on a rare event. Both the neural bridge and the guided proposal successfully capture the bimodal structure in the trajectories, as reflected by the two distinct clusters in the $X_{t,1}$ paths, particularly evident after $t\approx 0.5$. However, the guided proposal matches the reference distribution more closely, especially in the concentration and spread of trajectories beyond time $t=1.5$.
@@ 5.4 Stochastic Landmark Matching
Finally, we present a high-dimensional stochastic nonlinear conditioning task: stochastic landmark matching, see for instance Arnaudon et al. [2022]. We consider a stochastic model that describes the evolution of $n$ distinct landmarks in $\mathbb{R}^{d}$ of a closed nonintersecting curve in $\mathbb{R}^{d}$. The state at time $t$, $X_{t}$, consists of the concatenation of each of the $n$ landmark locations at time $t$. Hence, $X_{t}$ takes values in $\mathbb{R}^{dn}$.
The stochastic landmark model defined in Arnaudon et al. [2022], for ease of exposition considered without momentum variables, defines the process $(X_{t},\,t\in[0,T])$ as the solution to the SDE:
$\mathrm{d}X_{t}=Q(X_{t})\mathrm{d}W_{t},\quad Q(X_{t})_{ij}\coloneq k(X_{t}^{(i)},X_{t}^{(j)})\mathbf{I}_{d}.$ (32)
Here $\mathbf{I}_{d}$ is the $d$-dimensional identity matrix, $W$ is a $dn$-dimensional Wiener process and $k$ is a kernel function $\mathbb{R}^{d}\times\mathbb{R}^{d}\to\mathbb{R}$. In our numerical experiments we chose the Gaussian kernel $k(s,y)=\frac{1}{2}\alpha\exp\left(-\frac{\|x-y\|^{2}}{2\kappa^{2}}\right)$. The kernel parameters are chosen to be $\alpha=0.3,\kappa=0.5$ to ensure a strong correlation between a wide range of landmarks. Note that the diffusion coefficient $Q$ is state-dependent. To demonstrate the necessity of constructing such a state-dependent diffusion coefficient, consider the process $\tilde{X}$ solving the SDE
$\mathrm{d}\tilde{X}_{t}=Q(v)\mathrm{d}W_{t},$ (33)
where the diffusion constant is constant. As a result, the system becomes linear. The processes defined by Equation 32 and Equation 33 are fundamentally different. In particular, Equation 32 guarantees that $t\mapsto X_{t}$ generates a stochastic flow of diffeomorphisms [Sommer et al., 2021]. This diffeomorphic setting preserves the shape structure during evolution, whereas the linear process defined by Equation 33 does not. As illustrated in Figure 13, a visual comparison (using identical driving Wiener processes) reveals that the linear process disrupts the shape topology, leading to overlaps and intersections. The diffusion process defined by Equation 33 can however be used as auxiliary process in the construction of guided proposals. We opted for this choice in our numerical experiments.
In our numerical experiments, we chose one ellipse as the starting point and another ellipse as the endpoint of the bridge. Each ellipse is discretized by 50 landmarks, leading to the dimension of $X_{t}$ being 100. We took $T=1.0$. In the leftmost column of Figure 14 we fix a Wiener process and show on the top- and middle row the guided proposal and neural bridge using this Wiener process. We observed that due to the very simple choice of auxiliary process, the guided proposal has difficulty reaching the final state. In fact, we had to increase $\epsilon^{2}$ to $32e-$ 3 for not running into numerical instabilities. Here, one can see that the additional learning by the neural bridge gives much better performance. In the bottom row, we used the guided proposal with the same Wiener process as initialization, augmented by running $5000$ pCN iterations and plotted the final iteration. From this, one can see that these iterations provide another way to improve upon the guided proposal in the top row. The other columns repeat the same experiment with different Wiener process initializations. Due to the high-dimension of the problem, repeated simulation required for pCN steps may be computationally expensive.
5.5 Comparison of training time and number of parameters
We benchmarked our method on the four test cases described above, comparing it to two other deep learning-based approaches in Table 1. The comparison considers the number of network parameters and total training time, with all methods trained using the same number of gradient descent steps and batch sizes. Our method outperforms both alternatives in terms of model size and training efficiency.
OU ($d=1$) Cell ($d=2$) FHN ($d=2$) Landmark ($d=20$)
Methods #Params Time #Params Time #Params Time #Params Time
Adjoint forward [Baker et al., 2024b] $21,969$ $162.68s$ $22,114$ $265.06s$ N/A N/A $114,744$ $543.80s$
Score matching [Heng et al., 2022] $26,353$ $65.55s$ $26,498$ $103.38s$ N/A N/A $26,766$ $353.05s$
Neural guided bridge (ours) $921$ $44.12s$ $2,306$ $94.79s$ $3,362$ $113.77s$ $15,188$ $122.48s$
Table 1: Benchmarks with two other deep-learning based methods. Neither adjoint forward nor score matching is available for the FHN partially observed conditioning case (denoted as “N/A”, not applicable in the table). All the experiments are conducted with the parameters described in Appendix B.